### 20 Jan 2024 v1

# Handling incomplete outcomes and covariates in cluster-randomized trials: doubly-robust estimation, efficiency considerations, and sensitivity analysis

Abstract
In cluster-randomized trials, missing data can occur in various ways, including missing values in outcomes and baseline covariates at the individual or cluster level, or completely missing information for non-participants. Among the various types of missing data in CRTs, missing outcomes have attracted the most attention. However, no existing method comprehensively addresses all the aforementioned types of missing data simultaneously due to their complexity. This gap in methodology may lead to confusion and potential pitfalls in the analysis of CRTs. In this article, we propose a doubly-robust estimator for a variety of estimands that simultaneously handles missing outcomes under a missing-at-random assumption, missing covariates with the missing-indicator method (with no constraint on missing covariate distributions), and missing cluster-population sizes via a uniform sampling framework. Furthermore, we provide three approaches to improve precision by choosing the optimal weights for intracluster correlation, leveraging machine learning, and modeling the propensity score for treatment assignment. To evaluate the impact of violated missing data assumptions, we additionally propose a sensitivity analysis that measures when missing data alter the conclusion of treatment effect estimation. Simulation studies and data applications both show that our proposed method is valid and superior to the existing methods.
Keywords: cluster-randomized experiments, covariate adjustment, efficiency, machine learning, sensitivity.
1 Introduction
A cluster-randomized trial (CRT) is a type of randomized trial where the randomization units are clusters of individuals, such as classrooms, hospitals, or villages. Distinct from individually randomized trials, in a CRT, all individuals within the same cluster are assigned to the same treatment or control group, and the within-cluster observations are often correlated (Turner et al.,, 2017). In recent years, CRTs have been increasingly popular in evaluating treatment effects when the intervention is naturally designed to be implemented at the cluster level, when there is concern about within-cluster treatment contamination, and when administrative and policy considerations dictate cluster randomization.
Missing data is ubiquitous in CRTs. In a systematic review of CRTs published in 2011 (Díaz-Ordaz et al., 2014a, ), 95 out of 132 (72%) CRTs involved missing data, and the missing data proportion ranged 1%-47%. However, 66% CRTs adopted complete case analyses, which ignored all missing data. In a survey of 86 CRTs published between August 2013 and July 2014 (Fiero et al.,, 2016), 93% CRTs have missing outcomes with 55% of these trials using complete case analyses. Indeed, due to the multi-level data structure, missing data can occur in various ways in CRTs. At the individual level, a participant can have missing values in their outcome and covariates. At the cluster level, in addition to missing values among cluster-level covariates, a cluster may not have all its individuals enrolled, resulting in another type of missing data by design: no individual information is observed for non-participants; see Figure 1 for an example. In addition, since the cluster-population size (i.e., the total number of individuals in the cluster source population of interest) can also be unknown due to non-participation, more complexities arise in handling the joint missing data distribution in CRTs.
Figure 1: An illustration of missing data in a cluster of CRTs.
For addressing the missing data problem in CRTs, the majority of statistical literature focused on missing outcomes under the missing at random assumption. Multiple imputation (Hunsberger et al.,, 2001; Taljaard et al.,, 2008; Díaz-Ordaz et al., 2014b, ) is one of the earliest methods to model the missing outcome and covariates distribution in CRTs. As a model-based method, its validity relies on a correctly specified imputation model, which may be challenging to specify in practice (Turner et al.,, 2020). Prague et al., (2016) proposed an approach that combines augmented generalized estimating equations (Aug-GEE, Stephens et al.,, 2012) for outcome modeling and inverse probability weighting (IPW) to analyze CRTs with missing outcomes. A distinguishing feature of this estimator is its double-robustness, i.e., the resulting estimator is consistent as long as either the outcome model or the missing outcome propensity score is correctly specified, which offers additional robustness compared to multiple imputation. More recently, Balzer et al., (2023) used the targeted maximum likelihood estimation (TMLE) framework to model missing outcomes and derive doubly-robust estimators. Nugent et al., (2023) extended this method to handle within-cluster subsampling and missing values in a binary covariate; however, missing values in more general individual-level or cluster-level covariates were not considered. Furthermore, the existing TMLE framework implicitly assumed conditional independence among individuals (or among subclusters) within a cluster, which may not hold in all CRTs as the true within-cluster correlation structure is typically unknown. Simulation studies have also been extensively carried out to evaluate the performance of different methods; some of the most recent ones include Turner et al., (2020); Benitez et al., (2021); Kang et al., (2023). Despite these prior efforts, there has not been an approach that comprehensively addresses all kinds of missing data in CRTs as described in Figure 1.
In this work, our major contribution to CRTs is three-fold: first, we propose a new doubly-robust estimator to address missing data in outcomes, covariates, and cluster-population sizes; second, we offer three techniques to further improve precision and clarify when the semiparametric efficiency bound is attained; third, we contribute a novel sensitivity analysis for the missing data assumptions. In our proposed estimator, we adapt the IPW technique from Prague et al., (2016) to address missing outcomes and achieve double-robustness under a missing-at-random assumption. Missing covariates are imputed using the missingness-indicator method (Zhao and Ding,, 2022), which allows covariate misisng-not-at-random. For non-participation and missing cluster-population sizes, we provide a set of sufficient conditions—a uniform sampling framework—to achieve identification of our target estimand. In order to maximize the precision of the proposed estimator, we prove the statistical guarantee for each of the following techniques: choosing the optimal weight for individuals within a cluster, leveraging machine learning to learn the nuisance functions, and modeling the probability of treatment assignment. Each of the three techniques can be conveniently applied individually or in combination with others, offering practical tools to enhance the statistical power for analyzing CRTs. Finally, to accommodate potential violations of our missing data assumptions in CRTs, we propose a novel sensitivity analysis strategy to quantify the selection bias and the bias from outcome missing not at random. Specifically, by deriving the bias of our estimator with no missing data assumption, we compute the “tipping point”, which measures the difference between observed data and missing data distributions such that the null hypothesis cannot be rejected. The magnitude of the tipping point can then serve as a measure of sensitivity to violations of missing data assumptions. Taken together, we provide an all-in-one tool to simultaneously estimate the treatment effect estimand of interest, handle missing data in multiple domains, maximize precision, and offer a complementary sensitivity analysis framework to address key missing data assumptions.
The remainder of this article is organized as follows. We start with the simple setting with full enrollment in Section 2, where we formalize our framework and present the assumptions for identification of our estimands. In Section 3, we propose our estimator, derive their asymptotic results, and provide three approaches to further improve its precision. Section 4 generalizes the setting of Section 2 and results in Section 3 to accommodate within-cluster sampling and missing cluster-population sizes. In Section 5, we describe the new sensitivity analysis strategy for missing data assumptions. In Sections 6 and 7, we demonstrate our theoretical results via simulation studies and a real-data application. Section 8 concludes.
2 Notation and assumptions
We consider a CRT with $m$ clusters. Each cluster $i$ ($i=1,\dots,m$) contains $N_{i}$ individuals, while only $M_{i}$ individuals $(M_{i}\leq N_{i})$ are enrolled in the CRT. Here, $N_{i}$ is called the “cluster-population size” (i.e., the total number of individuals of interest), and $M_{i}$ is called the “sampled cluster size” (i.e., the number of individuals with observed data in the CRT). For each individual $j$ ($j=1,\dots,N_{i})$ in cluster $i$, we denote $Y_{ij}$ as the outcome of interest, $\boldsymbol{X}_{ij}=(X_{ij1},\dots,X_{ijp})^{\top}$ as a $p$-dimensional vector of individual-level baseline covariates, and $E_{ij}$ as a latent enrollment indicator such that $M_{i}=\sum_{j=1}^{N_{i}}E_{ij}$. At the cluster level, we further let $A_{i}$ represent a binary cluster-level treatment (with 1 for treatment and 0 for control) and $\boldsymbol{C}_{i}=(C_{i1},\dots,C_{iq})^{\top}$ be a $q$-dimensional vector of cluster-level baseline covariates.
In practice, outcomes, covariates, and cluster-population sizes are all subject to missingness. We define $R_{ij}^{Y}$ as a non-missing indicator of $Y_{ij}$ with $R_{ij}^{Y}=1$ if $Y_{ij}$ is observed and $R_{ij}^{Y}=0$ otherwise. Similarly, we define $R_{ijk}^{X},R_{il}^{C},R_{i}^{N}$ as the non-missing indicator for covariates $X_{ijk},C_{il},N_{i}$, respectively $(k=1,\dots,p;l=1,\dots,q)$. We assume that $A_{i}$ and $M_{i}$ are always observed by design, but the enrollment index set $\{j:E_{ij}=1\}$ is unobserved if $M_{i}<N_{i}$. To simplify the notation given missing covariates, we define $X_{ijk}^{*}=(R_{ijk}^{X},R_{ijk}^{X}X_{ijk})$ as the observed information for $X_{ijk}$; in the special case $R_{ijk}^{X}\equiv 1$, we set $X_{ijk}^{*}=X_{ijk}$. Following the same rule, we let $C_{il}^{*}=(R_{il}^{C},R_{il}^{C}C_{il})$ and $N_{i}^{*}=(R_{i}^{N},R_{i}^{N}N_{i})$ to denote the observed cluster-level covariate information and observed cluster-population size information. Finally, the observed data for each cluster become $\mathbf{O}_{i}=\{(R_{ij}^{Y},R_{ij}^{Y}Y_{ij},A_{i},\boldsymbol{B}_{ij}^{\top}):j=1,\dots,N_{i},E_{ij}=1\}$, where $\boldsymbol{B}_{ij}=(X_{ij1}^{*},\dots,X_{ijp}^{*},C_{i1}^{*},\dots,C_{iq}^{*},N_{i}^{*},M_{i})^{\top}$ is the observed baseline covariate data for individual $j$.
To begin with, we consider a basic setting where $M_{i}=N_{i}$, meaning that all individuals in a cluster are enrolled. In practice, this setting corresponds to scenarios where clusters represent household, class, or working group so that $N_{i}$ is relatively small. By focusing on this important basic setting, we are able to clearly present our results with key insights. In Section 4, we will return to the general setting where $M_{i}\leq N_{i}$ and provide sufficient conditions to generalize our results, and Section 5 will describe a sensitivity analysis in that general setting. Under this basic setting with $M_{i}=N_{i}$, the observed data simplifies to $\mathbf{O}_{i}=(\boldsymbol{R}_{i}^{Y},\boldsymbol{R}_{i}^{Y}\circ\boldsymbol{Y}_{i},A_{i},\mathbf{B}_{i})$, where $\boldsymbol{Y}_{i}=(Y_{i1},\dots,Y_{iN_{i}})^{\top}$, $\boldsymbol{R}_{i}^{Y}=(R_{i1}^{Y},\dots,R_{iN_{i}}^{Y})^{\top}$, and $\mathbf{B}_{i}=(\boldsymbol{B}_{i1},\dots,\boldsymbol{B}_{iN_{i}})^{\top}$ with $\circ$” denoting element-wise product between vectors. Furthermore, we define the outcome-complete data as $\boldsymbol{W}_{i}=(\boldsymbol{R}_{i}^{Y},\boldsymbol{Y}_{i},A_{i},\mathbf{B}_{i})$ and make the following structural assumptions on $\{\boldsymbol{W}_{1},\dots,\boldsymbol{W}_{m}\}$.
Assumption 1 (Super-population).
(1) $\{\boldsymbol{W}_{1},\dots,\boldsymbol{W}_{m}\}$ are independent and identically distributed samples from an unknown distribution $\mathcal{P}^{\boldsymbol{W}}$ with finite second moments. (2) Within each cluster $i$, the individual-level data $\{(R_{ij}^{Y},Y_{ij},\boldsymbol{B}_{ij}^{\top}):j=1,\dots,N_{i}\}$ are identically distributed given $(A_{i},N_{i})$.
Assumption 2 (Cluster randomization).
In the distribution $\mathcal{P}^{\boldsymbol{W}}$, $P(A_{i}=1)=\pi$ for some known constant $\pi\in(0,1)$, and $A_{i}$ is independent of $\mathbf{B}_{i}$.
Assumption 3 (Restricted missing at random, rMAR).
Each $R_{ij}^{Y}$ is conditionally independent of $\boldsymbol{Y}_{i}$ given $(A_{i},\mathbf{B}_{i})$ and $P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})$ is uniformly bounded away from 0.
Assumption 1(1) describes a super-population framework for inference and assumes inter-cluster independence, which is a standard setting in CRTs. While each $\boldsymbol{W}_{i}$ may have different dimensions due to varying $N_{i}$, they follow an identical distribution by a mixture data generating distribution $\mathcal{P}^{\boldsymbol{W}}=\mathcal{P}^{\boldsymbol{W}|N}\times\mathcal{P}^{N}$. Assumption 1(2) requires that, given the treatment condition and cluster-population size, each individual in a cluster is identically distributed, but they can be arbitrarily correlated. The motivation for this assumption is two-fold. First, while we label the individuals in a cluster from 1 to $N_{i}$ for notation convenience, they may not have a specific ordering in a cluster, and therefore it is reasonable to assume that one’s marginal distribution does not vary by their individual labels. Second, the assumption of marginal identical distribution facilitates the definition of our target estimands and the missing data mechanism, thereby simplifying the modeling and interpretation. Assumption 2 characterizes cluster randomization and assumes that missingness in covariates occurs prior to randomization. Beyond this mild condition, we make no further constraint on the covariate missingness mechanism; in other words, the baseline covariates can be missing not at random. Assumption 3, like the classical MAR assumption, induces conditional independence between missing-outcome indicators and outcomes given treatment and observed covariate data. Due to missingness in covariates, rMAR is stronger than MAR in that only observed data are conditioned on; that is, the missing outcome mechanism can depend on a covariate value $X_{ijk}$ if $R_{ijk}^{K}=1$ but not otherwise. In a non-clustered data setting, this type of assumption was also adopted by Rosenbaum and Rubin, (1984) and Chang et al., (2023) to handle missing covariates. Assumption 3 is also stronger than the censoring at random assumption for longitudinal studies, which assumes conditional independence between $Y_{ij}$ and $R_{ij}^{Y}$ given $A_{i},\mathbf{B}_{i}$ and all history outcomes under right censoring (Tsiatis,, 2006). This simplification is often necessary since there is no specific ordering of individuals within a cluster in the CRT setting and the concept of right censoring in longitudinal studies may not apply. While the rMAR assumption may not be verifiable from the observed data alone, in Section 5, we provide an approach to quantify the sensitivity of our proposed methods to violations of this rMAR assumption.
Given Assumptions 1-3, our goal is to estimate the average treatment effect, defined as
$\Delta=f(E[Y_{ij}|A_{i}=1],E[Y_{ij}|A_{i}=0]),$ (1)
where $f$ is a function defining the scale of effect measure, e.g., $f(x,y)=x-y$ for the difference scale and $f(x,y)=x(1-y)/\{(1-x)y\}$ for the odds ratio scale. Recently, increasing attention has been paid to estimands in cluster-randomized trials (Kahan et al.,, 2023), and the cluster-average treatment effect versus individual-average treatment effect are two typical estimands of interest. In short, the cluster-average treatment effect gives equal weight to each cluster and is interpreted as the treatment effect at the cluster level, while the individual-average treatment effect weighs each individual equally and refers to the treatment effect at the individual level. These two estimands are equal if $N_{i}$ is uncorrelated with $\boldsymbol{Y}_{i}$ in each arm $a$, which means the cluster-population size is non-informative. Here, Equation (1) defines the cluster-average treatment effect. If the individual-average treatment effect is of interest, we can replace $E[Y_{ij}|A_{i}=1]$ by $E[N_{i}]^{-1}E[N_{i}Y_{ij}|A_{i}=1]$, and our proposed methods below can be adapted following Wang et al., (2022) to target the individual-average treatment effect.
3 Doubly-robust estimation with incomplete outcome and covariate data in CRTs
3.1 Constructing the estimator
In the context of addressing missing outcomes in individually-randomized trials, doubly-robust estimation involves fitting two working models: an outcome regression model for observed outcomes and a propensity score model for outcome missingness. Based on the model fit, the doubly-robust estimator for the average treatment effect can be constructed in multiple ways (Robins et al.,, 2007; Rotnitzky et al.,, 2012). Under the MAR assumption, such estimators are all consistent if at least one of the two working models is correctly specified, thereby offering additional robustness to model misspecifications compared to conventional model-based inference.
For analyzing CRTs, we consider doubly-robust estimation in the presence of clustered data and estimate $\Delta$ with the following steps. First, we construct two parametric working models, an outcome regression model $\eta(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y})$ for $E[Y_{ij}|R_{ij}=1,A_{i},\mathbf{B}_{i}]$ and a propensity score function $\kappa(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{R})$ for $P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})$, where $\boldsymbol{\theta}_{Y}$ and $\boldsymbol{\theta}_{R}$ are nuisance parameters. For each individual $j$, both working models can depend on individual-specific information through $\boldsymbol{B}_{ij}$ and information from other individuals through $\mathbf{B}_{i}$. For example, we can set a linear outcome regression model $\eta(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y})=\beta_{0}+\beta_{A}A_{i}+\boldsymbol{\beta}_{\boldsymbol{B}}^{\top}\boldsymbol{B}_{ij}+\boldsymbol{\beta}_{\boldsymbol{X}^{o}}^{\top}\overline{\boldsymbol{X}_{i}^{o}}$ for $\overline{\boldsymbol{X}_{i}^{o}}$ being the cluster-average of observed individual-level covariates and $\boldsymbol{\theta}_{Y}=(\beta_{0},\beta_{A},\boldsymbol{\beta}_{\boldsymbol{B}}^{\top},\boldsymbol{\beta}_{\boldsymbol{X}^{o}}^{\top})^{\top}$. On top of this first-order model, within-cluster correlation structure can also be modeled via generalized estimating equations. Beyond this simple example, more complex models, such as penalized regression and spline regression, can also be used. In general, we do not restrict our proposed estimator to any specific working model, but only require $\boldsymbol{\theta}_{Y}$ and $\boldsymbol{\theta}_{R}$ are estimated by solving pre-specified estimating equations $\boldsymbol{\psi}_{Y}(\mathbf{O}_{i};\boldsymbol{\theta}_{Y})$ and $\boldsymbol{\psi}_{R}(\mathbf{O}_{i};\boldsymbol{\theta}_{R})$ so that $\sum_{i=1}^{m}\boldsymbol{\psi}_{Y}(\mathbf{O}_{i};\widehat{\boldsymbol{\theta}}_{Y})=\boldsymbol{0}$ and $\sum_{i=1}^{m}\boldsymbol{\psi}_{R}(\mathbf{O}_{i};\widehat{\boldsymbol{\theta}}_{R})=\boldsymbol{0}$. This setting covers a wide variety of estimators including maximum likelihood estimators.
In the next step, we define
$D_{ij}(a,\boldsymbol{\theta}_{Y},\boldsymbol{\theta}_{R})=\frac{w_{j}(N_{i})}{\sum_{j^{\prime}=1}^{N_{i}}w_{j^{\prime}}(N_{i})}\left[\frac{I\{A_{i}=a\}}{\pi^{a}(1-\pi)^{1-a}}\frac{R_{ij}^{Y}\{Y_{ij}-\eta(a,\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y})\}}{\kappa(a,\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{R})}+\eta(a,\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y})\right]$ (2)
for $a=0,1$ and arbitrary positive weights $w_{j}(N_{i})$. Then, we estimate $E[Y_{ij}|A_{i}=a]$ by $\widehat{\mu}_{a}=m^{-1}\sum_{i=1}^{m}\sum_{j=1}^{N_{i}}D_{ij}(a,\widehat{\boldsymbol{\theta}}_{Y},\widehat{\boldsymbol{\theta}}_{R})$ and compute $\widehat{\Delta}=f(\widehat{\mu}_{1},\widehat{\mu}_{0})$.
In the final step, we construct a variance estimator for $\widehat{\Delta}$ by sandwich variance estimation (Tsiatis,, 2006). Letting $\boldsymbol{\theta}=(\Delta,\mu_{1},\mu_{0},\boldsymbol{\theta}_{Y}^{\top},\boldsymbol{\theta}_{R}^{\top})^{\top}$ denote all parameters involved in the estimation procedure, the estimation of $\boldsymbol{\theta}$ in the first two steps can be equivalently expressed as $\sum_{i=1}^{m}\boldsymbol{\psi}(\mathbf{O}_{i};\widehat{\boldsymbol{\theta}})=\boldsymbol{0}$, where the joint estimating function is
$\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})=\left(\begin{array}[]{c}\Delta-f(\mu_{1},\mu_{0})\\ \sum_{j=1}^{N_{i}}D_{ij}(1,\boldsymbol{\theta}_{Y},\boldsymbol{\theta}_{R})-\mu_{1}\\ \sum_{j=1}^{N_{i}}D_{ij}(0,\boldsymbol{\theta}_{Y},\boldsymbol{\theta}_{R})-\mu_{0}\\ \boldsymbol{\psi}_{R}(\mathbf{O}_{i};\boldsymbol{\theta}_{R})\\ \boldsymbol{\psi}_{Y}(\mathbf{O}_{i};\boldsymbol{\theta}_{Y})\end{array}\right).$
The sandwich variance estimator $\widehat{Var}(\widehat{\Delta})$ is computed as the first-row first-column entry of
$\left\{\sum_{i=1}^{m}\frac{\partial}{\partial\boldsymbol{\theta}}\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})\bigg{|}_{\boldsymbol{\theta}=\widehat{\boldsymbol{\theta}}}\right\}^{-1}\left\{\sum_{i=1}^{m}\boldsymbol{\psi}(\mathbf{O}_{i};\widehat{\boldsymbol{\theta}})\boldsymbol{\psi}(\mathbf{O}_{i};\widehat{\boldsymbol{\theta}})^{\top}\right\}\left\{\sum_{i=1}^{m}\frac{\partial}{\partial\boldsymbol{\theta}}\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})\bigg{|}_{\boldsymbol{\theta}=\widehat{\boldsymbol{\theta}}}\right\}^{-1\ \top}.$
Our proposed estimator $\widehat{\Delta}$ reduces to the standard doubly-robust estimator for individual randomized trials if $N_{i}\equiv 1$. When $N_{i}>1$, the individual-level estimation formula is multiplied by a scaled weight $\{\sum_{j^{\prime}=1}^{N_{i}}w_{j^{\prime}}(N_{i})\}^{-1}w_{j}(N_{i})$ (Equation (2)) and summed up to obtain the cluster-level estimation formula. If the weight $w_{j}(N_{i})$ involves unknown parameters (e.g., the exchangeable correlation parameter $\rho$), we can estimate them as additional nuisance parameters and update the estimating functions $\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})$ and $\boldsymbol{\theta}$ to account for this change.
Our proposed estimator is related to but different from current estimators for CRTs. Without missing data, Equation (2) setting $w_{j}(N_{i})=1$ is equivalent to the targeted maximum likelihood estimation (Balzer et al.,, 2016) with parametric models for nuisance functions. Balzer et al., (2023) further proposed two-stage TMLE to handle missing outcomes, while weighting or missing covariates were not considered, and the propensity score models for missing outcomes were stratified by clusters. Prague et al., (2016) proposed a doubly-robust estimator based on GEE with a working model for the correlation matrix, assuming no missing covariates, $N_{i}$ is independent of $Y_{i}$ given $A_{i}$ (i.e., non-informative cluster-population sizes), and covariate adjustment is carried out in the form of Aug-GEE.
3.2 Double-robustness and asymptotic normality for the proposed estimator
To derive the asymptotic results for our proposed estimator, we assume the following regularity conditions on our estimating equations $\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})$.
Assumption 4 (regularity conditions for parametric nuisance models).
(1) $\boldsymbol{\theta}\in\mathbf{\Theta}$, a compact set in the Euclidean space. (2) $E[\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})]=\boldsymbol{0}$ has a unique solution $\boldsymbol{\theta}^{*}=(\Delta^{*},\mu_{1}^{*},\mu_{0}^{*},\boldsymbol{\theta}_{R}^{*},\boldsymbol{\theta}_{Y}^{*})$, an interior point of $\mathbf{\Theta}$. (3) $\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})$ has a finite second moment and second-order derivatives that are uniformly bounded by some integrable function $h(\mathbf{O}_{i})$ in a small neighborhood of $\boldsymbol{\theta}^{*}$. (4) $E[\frac{\partial}{\partial\boldsymbol{\theta}}\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{*}}]$ is invertible.
Assumption 4 consists of moment and continuity conditions to rule out degenerate conditions for the nuisance functions $\kappa$ and $\eta$, and does not make any parametric assumptions. Of note, Assumption 4(2) requires a unique root to $E[\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})]=\boldsymbol{0}$, which means a unique maximizer of the log-likelihood function in large samples if $\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})$ is the corresponding score function. Importantly, Assumption 4 is satisfied by most commonly used parametric models, including those mentioned in Section 3.1, as long as $\mathbf{O}_{i}$ has finite fourth moments. Given this additional assumption, the following theorem summarizes the consistency and asymptotic normality of our proposed estimator.
Theorem 1.
[…]
Theorem 1 provides the double-robustness property of our proposed estimator, which has two new implications. First, double-robustness is compatible with arbitrary intracluster correlation among $\boldsymbol{R}_{i}^{Y},\boldsymbol{Y}_{i}$, or $\mathbf{B}_{i}$ since modeling the intracluster correlation only changes the variance of our estimator but not consistency (and can be accounted for by the sandwich variance estimator). Second, covariate missing distributions have no impact on double-robustness under our assumptions, as long as we adjust for all observed covariates under the missingness indicator approach. This finding is consistent with the results of Zhao et al., (2023) on missing covariates in individually-randomized trials. In Section 3.3.5, we further discuss the influence of imputation methods for missing covariates. Together with the double-robustness, the asymptotic normality in Theorem 1 implies the large-sample validity of normal-approximated confidence intervals, hence providing a tool for hypothesis testing on $\Delta$ with asymptotically controlled Type I error.
3.3 Key considerations for improving precision
3.3.1 Choosing the optimal weights
For our proposed estimator, the weights $\boldsymbol{w}(N_{i})=\{w_{1}(N_{i}),\dots,w_{N_{i}}(N_{i})\}^{\top}$ corresponds to the intracluster correlation structure in generalized estimating equations. The conventional wisdom suggests an exchangeable correlation structure (i.e., $\boldsymbol{w}(N_{i})=\textup{Corr}(\rho)^{-1}\boldsymbol{1}_{N_{i}}$ with a correlation parameter $\rho$ and $\boldsymbol{1}_{N_{i}}$ being a $N_{i}$-dimensional vector of ones), while others suggest adopting constant weights, $\boldsymbol{w}(N_{i})=\boldsymbol{1}_{N_{i}}$, for improving robustness. Our derivation shows that, whichever weight we choose, it is important to offsets its impact on modifying the estimand by adding a factor $1/\boldsymbol{w}(N_{i})^{\top}\boldsymbol{1}_{N_{i}}$ such that we are not targeting $E[\boldsymbol{w}(N_{i})^{\top}\boldsymbol{1}_{N_{i}}]^{-1}E[\boldsymbol{w}(N_{i})^{\top}\boldsymbol{Y}_{i}|A_{i}=a]$, which can be different from $E[Y_{ij}|A_{i}=a]$ under informative cluster-population sizes. Once the weight is properly scaled, all weights will lead to consistent estimators, and, surprisingly, the exchangeable correlation structure $\boldsymbol{w}(N_{i})=\textup{Corr}(\rho)^{-1}\boldsymbol{1}_{N_{i}}$ will lead to the same estimator as independence correlation structure $\boldsymbol{w}(N_{i})=\boldsymbol{1}_{N_{i}}$ since $\textup{Corr}(\rho)^{-1}\boldsymbol{1}_{N_{i}}=\{1+(N_{i}-1)\rho\}^{-1}\boldsymbol{1}_{N_{i}}$ is proportional to $\boldsymbol{1}_{N_{i}}$. In addition to this interesting insight, we provide optimal weights in the following corollary.
Proposition 1.
Given the same conditions as in Theorem 1, we have
$\sqrt{m}(\widehat{\mu}_{a}-E[Y_{ij}|A_{i}=a])=\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\frac{\boldsymbol{w}(N_{i})^{\top}}{\boldsymbol{w}(N_{i})^{\top}\boldsymbol{1}_{N_{i}}}IF_{i}(a,\mathbf{O}_{i})+o_{p}(1),$
where $IF_{i}(a,\mathbf{O}_{i})$ is defined in the Supplementary Material for $a=0,1$. Then, the optimal weights $\boldsymbol{w}(N_{i})$ that minimizes the asymptotic variance of $\widehat{\Delta}$ is
$\boldsymbol{w}(N_{i})^{\textup{opt}}=c(N_{i})E\left[\left\{\sum_{a=0}^{1}f_{a}^{\prime}IF_{i}(a,\mathbf{O}_{i})\right\}\left\{\sum_{a=0}^{1}f_{a}^{\prime}IF_{i}(a,\mathbf{O}_{i})\right\}^{\top}\bigg{|}N_{i}\right]^{-1}\boldsymbol{1}_{N_{i}},$ (3)
where $(f_{1}^{\prime},f_{0}^{\prime})$ is the gradient of $f$ at $(E[Y_{ij}|A_{i}=1],E[Y_{ij}|A_{i}=0])$ and $c(N_{i})>0$ is an arbitrary scalar.
Proposition 1 sheds light on the recipe for improving the precision of the doubly-robust estimator by adjusting for the working correlation structure. To obtain the optimal weights, one needs to compute the matrix inverse in Equation (3) using observed data, which is often infeasible when we have a limited number of clusters.
However, under an important special case, $\boldsymbol{w}(N_{i})^{\textup{opt}}$ has a simple form. If we assume that the individual-level data in a cluster are exchangeable; that is, any permutation of their indices $j$ does not change their joint distribution. In practice, this is often the default setting widely used in trial planning. In this case, direct calculation shows that $\boldsymbol{w}(N_{i})^{\textup{opt}}$ is proportional to $\boldsymbol{1}_{N_{i}}$, indicating that the constant weight leads to optimal efficiency. Due to its optimality, we recommend using the simplest independence working correlation structure in practice, especially when the number of clusters is small and $N_{i}$ has large variability.
3.3.2 Estimating the nuisance functions using machine learning
An important question is how to specify the nuisance function estimators $\kappa(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\widehat{\boldsymbol{\theta}}_{R})$ and $\eta(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\widehat{\boldsymbol{\theta}}_{Y})$ such that the variance of $\widehat{\Delta}$ is minimized. For this purpose, Proposition 2 shows when correctly specified nuisance function estimators lead to the highest precision.
Proposition 2.
Given Assumptions 1-4 and pre-specified weights $\boldsymbol{w}(N_{i})$, if any of the following conditions holds:
  1. (1)
    $P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})$ is known,
  2. (2)
    or $P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})=\kappa(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{R}^{*})$ and $\boldsymbol{\theta}_{R}^{*}$ is consistently estimated using independent data source,
  3. (3)
    or $(R_{i1},\dots,R_{iN_{i}})$ are mutually independent given $(A_{i},\mathbf{B}_{i})$, and $P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})=\kappa(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{R}^{*})$,
then the estimator $\widehat{\Delta}$ is asymptotically most precise if $\eta(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y}^{*})=E[Y_{ij}|R_{ij}=1,A_{i},\mathbf{B}_{i}]$. The optimal asymptotic variance is
$v^{opt}=Var\left\{\sum_{a=0}^{1}\sum_{j=1}^{N_{i}}f_{a}^{\prime}\frac{w_{j}(N_{i})}{\sum_{j^{\prime}=1}^{N_{i}}w_{j^{\prime}}(N_{i})}U_{ij}(a,\mathbf{O}_{i})\right\},$ (4)
where $(f_{1}^{\prime},f_{0}^{\prime})$ is the gradient of $f$ at $(E[Y_{ij}|A_{i}=1],E[Y_{ij}|A_{i}=0])$ and
$U_{ij}(a,\mathbf{O}_{i})=\frac{I\{A_{i}=a\}}{\pi^{a}(1-\pi)^{1-a}}\frac{R_{ij}^{Y}\{Y_{ij}-E[Y_{ij}|R_{ij}=1,A_{i}=a,\mathbf{B}_{i}]\}}{P(R_{ij}^{Y}=1|A_{i}=a,\mathbf{B}_{i})}+E[Y_{ij}|R_{ij}=1,A_{i}=a,\mathbf{B}_{i}].$
Proposition 2 lists three scenarios where, given either known or correctly specified models for the missing outcome propensity score, a correctly specified outcome model $\eta$ yields the most precise estimator asymptotically. Among the three conditions, condition (1) is the most restrictive one since it requires a known mechanism for missing outcomes. In practice, this condition is plausible if outcomes are missing completely at random or if outcome missingness only depends on some discrete variables. Condition (2) is weaker and assumes a correctly specified missing outcome model, but requires the missing outcome model to be estimated using independent data source, such as historical data or cross-fitting (which we introduce later in this subsection). If, otherwise, we use the same data to consistently estimate $\boldsymbol{\theta}_{R}^{*}$ and $\Delta$, then the estimation of $\boldsymbol{\theta}_{R}^{*}$ will result in an extra term (i.e., a projection term to the nuisance tangent space) in the influence function of $\widehat{\Delta}$, as implied by the semiparametric theory (Tsiatis,, 2006). Because of this extra term, the asymptotic variance of $\widehat{\Delta}$ also involves $E[R_{ij}R_{ij^{\prime}}|A_{i},\mathbf{B}_{i}]$ for $j\neq j^{\prime}$, which may cause the correctly specified outcome model to be suboptimal in certain data-generating distributions. However, if we consider a special scenario where $(R_{i1},\dots,R_{iN_{i}})$ are mutually independent given $(A_{i},\mathbf{B}_{i})$, as stated in condition (3), then the correctly specified outcome model is optimal. Overall, Proposition 2 is a generalization of Tan, (2007); Tsiatis and Davidian, (2007); Cao et al., (2009) from independent individual data to CRTs.
To maximize the precision of $\widehat{\Delta}$, Proposition 2 motivates us to find nuisance function estimators that contain the truth and, ideally, are constructed with independent data sources. However, parametric models only search for the truth in a relatively small space, and are often too restrictive to capture the truth, e.g., non-linear functions. Furthermore, independent data sources are rarely available, especially in the CRT setting. To overcome these challenges, we propose to use machine learning with cross-fitting for nuisance function estimation to improve precision. This procedure has the following steps. First, we randomly partition $m$ clusters into $K$ parts such that the numbers of clusters across parts differ by at most 1. We denote the observed data for part $k$ as $\mathcal{Q}_{k}$ and the rest as $\mathcal{Q}_{-k}$. Next, for each $k=1,\dots,K$, we use data $\mathcal{Q}_{-k}$ to train machine learning models $\widehat{\kappa}^{(k)}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})$ and $\widehat{\eta}^{(k)}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})$ for $P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})$ and $E[Y_{ij}|R_{ij}=1,A_{i},\mathbf{B}_{i}]$, respectively, and evaluate the trained models on $\mathcal{Q}_{k}$. Finally, the nuisance function estimators $\widehat{\kappa}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})$ and $\widehat{\eta}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})$ are the combinations of $\widehat{\kappa}^{(k)}$ and $\widehat{\eta}^{(k)}$ evaluated at $\mathcal{Q}_{k}$ for all $k$. In practice, we recommend choosing $K$ such that $m/K\geq 10$. For the machine learning models, we can choose any method such that the following regularity condition holds.
Assumption 5 (Regularity conditions for machine learning models).
(1) $\widehat{\kappa}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})-P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})=o_{p}(m^{-1/4})$ and $\widehat{\eta}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})-E[Y_{ij}|R_{ij}=1,A_{i},\mathbf{B}_{i}]=o_{p}(m^{-1/4})$ in $L_{2}$-norm. (2) $\widehat{\kappa}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})$ is uniformly bounded in $(0,1)$. (3) $N_{i}$ is upper bounded by some large constant.
Assumption 5(1) places constraints on the convergence rate of the nuisance function estimators to the truth to be faster than $m^{-1/4}$. This condition is much weaker than the classical $m^{-1/2}$ rate and can be achieved by many machine learning algorithms, such as deep neural networks (Farrell et al.,, 2021), random forests (Wager and Walther,, 2015), and highly-adaptive lasso (Benkeser and Van Der Laan,, 2016). For our setting of clustered data, we can simplify the estimator by setting $\widehat{\kappa}(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i})=\widehat{\kappa}(A_{i},\boldsymbol{B}_{ij},h(\mathbf{B}_{i}))$ for a summary, potentially high-dimensional function $h$ such that the above machine learning models can be directly applied. Emmenegger and Bühlmann, (2023) also provided tools to combine machine learning with mixed-effect models for handling clustered data. In addition, we assume Assumptions 5(2) and (3) to control the remainder terms and estimate variance.
With the new nuisance function estimators, we update Equation (2) correspondingly and estimate $\widehat{\Delta}^{\textup{ml}}$ in the same way. Its variance estimator, $\widehat{Var}(\widehat{\Delta}^{\textup{ml}})$, is constructed using cross-fitting and provided in the Supplementary Material. Theorem 2 below characterizes the asymptotic behavior of this estimator.
Theorem 2.
[…]
Theorem 2 differs from Theorem 1 in that it achieves consistency and the optimal variance, borrowing the power from machine learning. Compared to $\widehat{\Delta}$ using potentially misspecified parameter nuisance function estimators, $\widehat{\Delta}^{\textup{ml}}$ has higher asymptotic precision, which can be translated into narrower confidence intervals and fewer clusters to achieve a desired statistical power. However, machine-learning models often require a moderate to large number of clusters to avoid over-fitting. Furthermore, the cross-fitting procedure may also be prone to finite-sample bias especially when the sample size is limited. In contrast, parsimonious parametric models, while restrictive, have the advantage in stability and may often outperform machine learning models given few clusters. Therefore, we recommend considering machine learning to optimize variance when the number of clusters is large, say, $m\geq 80$.
3.3.3 Efficiency bounds
We have just derived the optimal variance for our proposed estimator $\widehat{\Delta}$ with machine learning, but is there an estimator, potentially with a different form, that is more precise than $\widehat{\Delta}^{\textup{ml}}$? To answer this question, we need to derive the efficiency bounds for $\Delta$ under our framework. To start with, we present a special case where $\widehat{\Delta}^{\textup{ml}}$ is efficient.
Proposition 3.
Given Assumptions 1-3, and assuming that $\{(Y_{ij},R_{ij}):j=1,\dots,N_{i}\}$ are mutually independent given $(A_{i},\mathbf{B}_{i})$, then $v^{opt}$ defined in Equation (4) with $w_{j}(N_{i})=1$ is the asymptotic variance lower bound for any regular and asymptotic linear estimator.
Proposition 3 implies that, when all the intracluster correlations among outcomes and missing outcome indicators are caused by observed covariates, $\widehat{\Delta}^{\textup{ml}}$ (and $\widehat{\Delta}$ with both nuisance function models correctly specified) achieves the efficiency lower bound asymptotically.
In the general case, however, $v^{opt}$ may not be the efficiency lower bound, but the true lower bound may not be worth pursuing in the CRT setting. Since $\{(Y_{ij},R_{ij}):j=1,\dots,N_{i}\}$ can be correlated, achieving the efficiency bound requires modeling $P(\boldsymbol{R}_{i}^{Y}=\boldsymbol{r}|A_{i},\mathbf{B}_{i})$ for any $N_{i}$-dimensional binary vector $\boldsymbol{r}$ in the support of $\boldsymbol{R}_{i}^{Y}$, and modeling $E[Y_{ij}|\boldsymbol{r}\circ\boldsymbol{Y}_{i},A_{i},\mathbf{B}_{i}]$, i.e., the conditional mean of unobserved outcomes given all observed outcomes, treatment, and observed covariates. (See Theorem 7.2 of Tsiatis, (2006) for technical details.) Intuitively, these two models utilize more information than $P(R_{ij}^{Y}=1|A_{i},\mathbf{B}_{i})$ and $E[Y_{ij}|R_{ij}=1,A_{i},\mathbf{B}_{i}]$ to predict $R_{ij}^{Y}$ and $Y_{ij}$, thereby further improving precision. However, when $N_{i}$ is large, as in CRTs, these two distributions can take up to $2^{N_{i}}$ unique functions, which are often too many to be accurately learned from observed data. Therefore, while we acknowledge the existence of a more efficient estimator than $\widehat{\Delta}$ and $\widehat{\Delta}^{\textup{ml}}$ asymptotically, such an estimator is often infeasible to obtain in finite samples and impractical, and, therefore, we do not further pursue such an estimator.
3.3.4 Modeling the propensity score for treatment assignment
In individually randomized trials, modeling the propensity score for treatment assignment has been established as a technique to improve precision (Shen et al.,, 2014; Zeng et al.,, 2021). While the distribution of $A_{i}$ given covariates is known to be $\pi$, constructing an estimator for it (and using the same data to estimate $\Delta$) will result in an extra term in the influence function of $\widehat{\Delta}$, and this extra term can be equivalently obtained by fitting an outcome regression model, which explains why this seemingly unnecessary model for treatment assignment can improve precision.
Motivated by this idea, we also consider modeling $P(A_{i}=1|\mathbf{B}_{i})=\pi(\mathbf{B}_{i};\boldsymbol{\theta}_{A})$ via maximum likelihood estimation, and assume that $\pi(\mathbf{B}_{i};\boldsymbol{\theta}_{A})=\pi$ for some $\boldsymbol{\theta}_{A}$. The estimator $\widehat{\mu}_{a}$ is then updated by substituting $\pi(\mathbf{B}_{i};\widehat{\boldsymbol{\theta}}_{A})$ for $\pi$, and the estimating functions $\boldsymbol{\psi}(\mathbf{O}_{i};\boldsymbol{\theta})$ is updated to account for the estimation of $\boldsymbol{\theta}_{A}$. With this change, the following proposition characterizes the benefits of modeling treatment assignments.
Proposition 4.
Assume Assumptions 1-3 and a correctly specified model $\kappa(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{R})$ for outcome missingness [^1] In the sense that the estimating function for $\boldsymbol{\theta}_{R}$ has mean zero conditioning on $(A_{i},\mathbf{B}_{i})$, which automatically holds under generalized estimating equations or maximum likelihood estimation assuming conditional independence. For arbitrary methods to estimate $\boldsymbol{\theta}_{R}$, this extra assumption rules out estimating functions that involve artificial terms, like $(A-\pi)\mathbf{B}_{i}$, beyond functions of $\boldsymbol{\theta}_{R}$ that is needed for parameter estimation.. Then, given any weight $\boldsymbol{w}(N_{i})$ and any outcome model $\eta(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y})$, modeling $\pi(\mathbf{B}_{i};\boldsymbol{\theta}_{A})$ does not increase the asymptotic variance of $\widehat{\Delta}$, and it does not change the asymptotic variance of $\widehat{\Delta}$ when the outcome model $\eta(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y})$ is correctly specified.
Proposition 4 generalizes the results for inverse-probability weighted estimator from individually randomized trials to CRTs with missing data (Chang et al.,, 2023; Zhao et al.,, 2023). When the model for outcome missingness is correctly specified, adding a model for treatment assignment guarantees no efficiency loss asymptotically, which justifies the benefits of this technique in CRTs. For example, when $\eta(A_{i},\boldsymbol{B}_{ij},\mathbf{B}_{i};\boldsymbol{\theta}_{Y})$ is set to be zero, $\widehat{\Delta}$ reduces to the inverse-probability weighted estimator, and modeling $\pi(\mathbf{B}_{i};\boldsymbol{\theta}_{A})$ can improve precision as long as the covariates are prognostic. When there are no or few missing outcomes, modeling $\pi(\mathbf{B}_{i};\boldsymbol{\theta}_{A})$ can also supplement the outcome model to improve precision. For constructing the model $\pi(\mathbf{B}_{i};\boldsymbol{\theta}_{A})$, one can use a logistic regression model on intercept and cluster-summary variables, which ensures that $\pi(\mathbf{B}_{i};\boldsymbol{\theta}_{A})$ is correctly specified.
3.3.5 Handling missing covariates
Recall the definition of $X_{ijk}^{*}=(R_{ijk}^{X},R_{ijk}^{X}X_{ijk})$. At first glance, the product $R_{ijk}^{X}X_{ijk}$ imputes all missing covariates by zero. However, $X_{ijk}^{*}$ in fact encompasses a wide range of imputation methods because the definition of $X_{ijk}^{*}$ captures all observed information. For example, if missing values in $X_{ijk}$ are imputed by a constant $c$, then the imputed covariate is $R_{ijk}^{X}X_{ijk}+(1-R_{ijk}^{X})c$, which is a function of $X_{ijk}^{*}$. More generally, if imputation for missing covariates is a function of other observed covariates, then the imputed covariate vector remains a function of $\boldsymbol{B}_{ij}$. Since our proposed method is not restricted to a specific function form of $\boldsymbol{B}_{ij}$ for nuisance function estimation, our proposed estimator is compatible with many covariate-imputation methods. Specifically, letting $\widetilde{\boldsymbol{B}}_{ij}=h(\boldsymbol{B}_{ij};\boldsymbol{\theta}_{B})$ denote the imputed covariate data under an imputation model $h$ with parameters $\boldsymbol{\theta}_{B}$, the nuisance function models based on imputed covariate data can be $\widetilde{\kappa}(a,\widetilde{\boldsymbol{B}}_{ij};\widetilde{\boldsymbol{\theta}}_{R})$ and $\widetilde{\eta}(a,\widetilde{\boldsymbol{B}}_{ij};\widetilde{\boldsymbol{\theta}}_{Y})$. This two-step procedure corresponds to specifying $\kappa(a,\boldsymbol{B}_{ij};\boldsymbol{\theta}_{R})=\widetilde{\kappa}(a,h(\boldsymbol{B}_{ij};\boldsymbol{\theta}_{B});\widetilde{\boldsymbol{\theta}}_{R})$ and $\eta(a,\boldsymbol{B}_{ij};\boldsymbol{\theta}_{Y})=\widetilde{\eta}(a,h(\boldsymbol{B}_{ij};\boldsymbol{\theta}_{B});\widetilde{\boldsymbol{\theta}}_{R})$, and it therefore can be viewed as a choice of nuisance function modeling on $\boldsymbol{B}_{ij}$. When there is no missing outcome, the choice of these covariate-imputation methods has no impact on the validity of the proposed estimator asymptotically, since it is robust to any misspecification of the outcome regression model. Given missing outcomes, we need at least one of the two nuisance functions to be correctly specified such that consistency is achieved.
When a generalized linear model is used to linearly adjust for $X_{ijk}^{*}$, then constant imputation for missing covariates is equivalent to zero imputation. This is because $\beta_{0}+\beta_{1}R_{ijk}^{X}+\beta_{2}R_{ijk}^{X}X_{ijk}=(\beta_{0}-\beta_{2}c)+(\beta_{1}+\beta_{2}c)R_{ijk}^{X}+\beta_{2}\{R_{ijk}^{X}X_{ijk}+(1-R_{ijk}^{X})c\}$ for arbitrary parameters $(\beta_{0},\beta_{1},\beta_{2})$, meaning that the model specification is mathematically the same regardless of the imputation value $c$. Therefore, it has been recommended to simply use zero imputation for such model specifications (Zhao and Ding,, 2022; Zhao et al.,, 2023). For nonlinear models, however, different imputation methods may result in changes in nuisance models, and their impact on efficiency can be intractable and may require a case-by-case examination.
4 Estimation under uniform within-cluster sampling
When a cluster only enrolls a proportion of individuals into the CRT, we have $M_{i}<N_{i}$ and $N_{i}$ is subject to missingness. This setting often corresponds to CRTs with large clusters such as villages, cities, or health systems, where we aim to generalize the treatment effect estimated from the $M_{i}$ samples to all $N_{i}$ individuals. However, since neither outcomes nor individual-level covariates are observed for non-participants, the enrollment mechanism (i.e., $P(E_{ij}=1|Y_{ij},\boldsymbol{B}_{ij})$) is unidentifiable from observed data. In this case, we provide sufficient conditions, Assumption 6 in place of Assumption 3, to characterize a basic scenario so that doubly-robust estimators constructed for $\Delta$ target the estimand of interest.
Assumption 6 (rMAR with uniform sampling).
(1)For each $i=1,\dots,m$, the enrollment vector $\boldsymbol{E}_{i}=(E_{i1},\dots,E_{iN_{i}})$ is independently generated by $P(\boldsymbol{E}_{i}=\boldsymbol{e}|M_{i},N_{i})={\binom{N_{i}}{M_{i}}}^{-1}$ for any $N_{i}$-dimensional binary vector $\boldsymbol{e}$ with $M_{i}$ ones. (2) Given observed cluster-level covariates $(C_{i1}^{*},\dots,C_{iq}^{*},M_{i})$ and $R_{i}^{N}=0$, $N_{i}$ is conditionally independent of $(R_{ij}^{Y},Y_{ij},X_{ij1}^{*},\dots,X_{ijp}^{*})$. (3) $R_{ij}^{Y}$ is conditionally independent of $Y_{ij}$ given $(E_{ij}=1,A_{i},\boldsymbol{B}_{ij})$ for each $j$ with $E_{ij}=1$.
Assumption  6(1) characterizes the uniform distribution of $\boldsymbol{E}_{i}$. Given the sampled cluster size and cluster-population size, it assumes that $\boldsymbol{E}_{i}$ is independent of all other information (including covariates, treatment assignment, and outcomes), making $\boldsymbol{E}_{i}$ causing no selection bias. If, otherwise, $E_{ij}$ is correlated with $Y_{ij}$ or $\boldsymbol{X}_{ij}$ given $(M_{i},N_{i})$, then the treatment effect can be different between participants and non-participants, for which we discuss the identifiability of $\Delta$ in Section 5. Assumption 6(2) handles missing $N_{i}$ by assuming non-informative cluster-population sizes. In other words, cluster-level covariates explain all the correlations between $N_{i}$ and individual-level data whenever $N_{i}$ is unobserved. This assumption can be reasonable if we observe a proxy variable for $N_{i}$ (e.g., a categorical variable encoding the level of cluster-population sizes by small, medium, and large) and the individual data distribution is invariant to small perturbations of $N_{i}$. Finally, Assumption 6(3) updates Assumption 3 to accommodate sampling and changes in observed data. Here, an rMAR mechanism is still assumed, while the conditioning variables are restricted to $E_{ij}=1$ (such that the data for individual $j$ can be observed), $A_{i}$, and $\boldsymbol{B}_{ij}$ (instead of $\mathbf{B}_{i}$ in Assumption 3, which involves covariate data from non-participants). Altogether, Assumption 6 provides a scenario where the trial participants are representative samples of the entire cluster and missing data are ignorable given observed data.
Under this setting, we specify the parametric nuisance function estimators as $\eta(A_{i},\boldsymbol{B}_{ij};\boldsymbol{\theta}_{Y})$ and $\kappa(A_{i},\boldsymbol{B}_{ij};\boldsymbol{\theta}_{R})$ and compute $\widehat{\boldsymbol{\theta}}_{Y}$ and $\widehat{\boldsymbol{\theta}}_{R}$ in the same way as in Section 3.1. We next define
$D_{ij}^{\dagger}(a,\boldsymbol{\theta}_{Y},\boldsymbol{\theta}_{R})=\frac{E_{ij}}{M_{i}}\left[\frac{I\{A_{i}=a\}}{\pi^{a}(1-\pi)^{1-a}}\frac{R_{ij}^{Y}\{Y_{ij}-\eta(a,\boldsymbol{B}_{ij};\boldsymbol{\theta}_{Y})\}}{\kappa(a,\boldsymbol{B}_{ij};\boldsymbol{\theta}_{R})}+\eta(a,\boldsymbol{B}_{ij};\boldsymbol{\theta}_{Y})\right]$
for $j=1,\dots,N_{i}$ and $a=0,1$. The estimator for $\Delta$ is then computed as $\widehat{\Delta}^{\dagger}=f(\widehat{\mu}_{1}^{\dagger},\widehat{\mu}_{0}^{\dagger})$, where $\widehat{\mu}_{a}^{\dagger}=m^{-1}\sum_{i=1}^{m}\sum_{j=1}^{N_{i}}D_{ij}^{\dagger}(a,\widehat{\boldsymbol{\theta}}_{Y},\widehat{\boldsymbol{\theta}}_{R})$ for $a=0,1$. For constructing the variance estimator $\widehat{Var}(\widehat{\Delta}^{\dagger})$, we again use the sandwich variance estimation as described in Section 3.1. Compared to $\widehat{\Delta}$ defined without within-cluster sampling, the major difference of $\widehat{\Delta}^{\dagger}$ is the weight $M_{i}^{-1}E_{ij}$ in place of $\{\sum_{j^{\prime}=1}^{N_{i}}w_{j^{\prime}}(N_{i})\}^{-1}w_{j}(N_{i})$, representing constant weights among sampled individuals.
For deriving the asymptotic results, we assume the same regularity conditions as in Assumption 4 for $\widehat{\Delta}^{\dagger}$, but do not repeat here for brevity. Theorem 3 below provides parallel results of double-robustness and asymptotic normality for $\widehat{\Delta}^{\dagger}$ to Theorem 1.
Theorem 3.
[…]
Beyond parametric models, machine learning can be incorporated into $\widehat{\Delta}^{\dagger}$ in the same way as $\widehat{\Delta}^{\textup{ml}}$, and Propositions 2-4 apply under an additional assumption that $(R_{ij}^{Y},Y_{ij})$ is conditionally independent of $\boldsymbol{B}_{ij^{\prime}}$ given $(A_{i},\boldsymbol{B}_{ij})$ for $j\neq j^{\prime}$; this additional assumption is needed to accommodate the change in the rMAR assumption and nuisance function estimators. For completeness, we provide the details in the Supplementary Material.
5 Sensitivity analyses for missing data assumptions
In order to estimate $\Delta$, we have made structural assumptions to deal with the sampling procedure and missingness in outcomes. However, these assumptions may not hold in practice. For example, enrolled individuals may present a larger treatment effect because they represent specific population subgroups. Among the enrolled individuals, outcomes can also be missing not at random; that is, missingness depends on outcome values and unobserved covariates. To evaluate such effects, we study the bias of our proposed estimators without putting any constraint on the sampling or missing data mechanism. To simplify the discussion and present a relatively clean result, we assume $N_{i}$ is observed, $E_{ij}$ is a pre-randomization variable, and all individual-level data in a cluster are identically distributed; these assumptions are formally presented in Proposition 5 below.
Proposition 5.
Assume
  1. (1)
    $\widetilde{\boldsymbol{W}}_{i}=\{(R_{ij}^{Y},Y_{ij},A_{i},E_{ij},\boldsymbol{B}_{ij}),j=1,\dots,N_{i}\}$ for $i=1,\dots,m$ are independent and identically distributed. Within $\widetilde{\boldsymbol{W}}_{i}$, the individual data $(R_{ij}^{Y},Y_{ij},E_{ij},\boldsymbol{B}_{ij})$ for each $j$ are also identically distributed given $A_{i},N_{i}$ (where $N_{i}$ is assumed to be observed),
  2. (2)
    $A_{i}$ is independent of $(E_{ij},\boldsymbol{B}_{ij})$.
Given the Regularity conditions in Assumption 4, then $\widehat{\mu}_{a}^{\dagger}$ defined in Section 4 satisfies
[…]
if either nuisance function model $\kappa$ or $\eta$ is correctly specified. Furthermore,
$\underline{\mu}_{a}-E[Y_{ij}|A_{i}=a]=E\left[\frac{N_{i}-M_{i}}{N_{i}}\delta_{a}(M_{i},N_{i})\right]+E[E[(1-R_{ij}^{Y})\gamma_{a}(\boldsymbol{B}_{ij})|E_{ij}=1,A_{i}=a,M_{i},N_{i}]],$ (5)
where
$\delta_{a}(M_{i},N_{i})$ $=E[Y_{ij}|E_{ij}=1,A_{i}=a,M_{i},N_{i}]-E[Y_{ij}|E_{ij}=0,A_{i}=a,M_{i},N_{i}],$
$\gamma_{a}(\boldsymbol{B}_{ij})$ $=E[Y_{ij}|R_{ij}^{Y}=1,E_{ij}=1,A_{i}=a,\boldsymbol{B}_{ij}]-E[Y_{ij}|R_{ij}^{Y}=0,E_{ij}=1,A_{i}=a,\boldsymbol{B}_{ij}].$
Equation (5) quantifies the asymptotic bias of $\widehat{\mu}_{a}^{\dagger}$ in estimating $E[Y_{ij}|A_{i}=a]$. The bias has two components, the first characterizing the selection bias from sampling and the second showing the bias from outcome missing not at random. For the first component, $E\left[\frac{N_{i}-M_{i}}{N_{i}}\delta_{a}(M_{i},N_{i})\right]$ is determined by the outcome mean difference between participants and non-participants, $\delta_{a}(M_{i},N_{i})$, and the proportion of non-participants, $\frac{N_{i}-M_{i}}{N_{i}}$. This source of bias increases if either $\delta_{a}(M_{i},N_{i})$ or $\frac{N_{i}-M_{i}}{N_{i}}$ becomes larger. Similarly, the second component, $E[E[(1-R_{ij}^{Y})\gamma_{a}(\boldsymbol{B}_{ij})|E_{ij}=1,A_{i}=a,M_{i},N_{i}]]$, is determined by the conditional mean difference between missing outcomes and observed outcomes among participants, $\gamma_{a}(\boldsymbol{B}_{ij})$, and the conditional proportion of missing outcomes, $P(R_{ij}^{Y}=0|E_{ij}=1,A_{i}=a,\boldsymbol{B}_{ij})$. Roughly speaking, when more outcomes are missing, and missing outcomes have larger difference from observed outcomes, the bias of $\underline{\mu}_{a}$ becomes larger. Finally, when all individuals are enrolled as considered in Section 3, Equation (5) still holds for $\widehat{\mu}_{a}$ with the first component of bias being zero.
While Equation (5) involves inestimable quantities, it can still be used to perform a sensitivity analysis. The sensitivity analysis aims to identify the “tipping point”, i.e., increasing $|\delta_{a}(M_{i},N_{i})|$ and $|\gamma_{a}(\boldsymbol{B}_{ij})|$ such that $\widehat{\Delta}$ is no longer statistically significant. (If $\widehat{\Delta}$ is insignificant at the beginning, the sensitivity analysis can visualize the impact of missing data on bias and provide qualitative evidence.) The larger the tipping point is, the less sensitive our proposed estimator is to violations of missing data assumptions. For facilitating the practical implementation, one can set $\delta_{a}(M_{i},N_{i})\equiv\delta_{a}$ and $\gamma_{a}(\boldsymbol{B}_{ij})\equiv\gamma_{a}$ to find constant tipping points. As an example, when $\Delta=E[Y_{ij}|A_{i}=1]-E[Y_{ij}|A_{i}=0]$, then the asymptotic bias of $\widehat{\Delta}^{\dagger}$ is
$E\left[\frac{N_{i}-M_{i}}{N_{i}}\right](\delta_{1}-\delta_{0})$
$\quad+E[P(R_{ij}^{Y}=0|A_{i}=1,E_{ij}=1,M_{i},N_{i})]\gamma_{1}-E[P(R_{ij}^{Y}=0|A_{i}=0,E_{ij}=1,M_{i},N_{i})]\gamma_{0},$
where $E\left[\frac{N_{i}-M_{i}}{N_{i}}\right]$ and $P(R_{ij}^{Y}=0|A_{i}=a,E_{ij}=1,M_{i},N_{i})$ are estimable from data. Then a grid search of tipping points on $(\delta_{1}-\delta_{0},\gamma_{1},\gamma_{0})$ can reflect the sensitivity of our estimator. Even when $N_{i}$ is unobserved, we can use bounds for $E\left[\frac{N_{i}-M_{i}}{N_{i}}\right]$ and approximations of $P(R_{ij}^{Y}=0|A_{i}=a,E_{ij}=1,M_{i},N_{i})$ by $P(R_{ij}^{Y}=0|A_{i}=a,E_{ij}=1)$ to perform the sensitivity analysis. In the special case with $N_{i}=1$, this approach reduces to the pattern mixture model for individual randomized trials (Little,, 1993; Little et al.,, 2012).
Beyond this sensitivity analysis, Equation (5) can also be used for bias correction under certain missing-not-at-random assumptions. For example, one potentially conservative method (Gao et al.,, 2023) is to assume $E[Y_{ij}|R_{ij}^{Y}=0,E_{ij}=1,A_{i}=a,\boldsymbol{B}_{ij}]=E[Y_{ij}|R_{ij}^{Y}=1,E_{ij}=1,A_{i}=0,\boldsymbol{B}_{ij}]$ for $a=0,1$, meaning that the missing outcomes have the same mean as if they are observed in the control arm. As a result, $\gamma_{a}(\boldsymbol{B}_{ij})$ is estimable for observed data, and the bias term from outcome missing not at random can be fully corrected. Following a similar assumption for $E_{ij}$, we can correct the selection bias and achieve doubly-robust identification.
6 Simulation study
We conduct two simulation studies, the first confirming our theoretical results in Sections 3 and 4 and the second demonstrating the sensitivity analysis method proposed in Section 5.
6.1 Simulation design
In the first simulation study, we consider the combination of eight scenarios defined by (i) no sampling (under Assumption 3) versus uniform sampling with missing $N_{i}$ (under Assumption 6), (ii) a small ($m=30$) versus large ($m=100$) number of clusters, and (iii) a small ($p_{m}=10\%$) versus moderate ($p_{m}=30\%$) proportion of missing data for both outcomes and covariates.
For each scenario, we independently generate data for each cluster $i$ as follows. First, we sample the cluster-population size $N_{i}$ from a uniform distribution on integers between 10 and 90, a cluster-level covariate $C_{i}\sim N(N_{i}/10,1)$ with its non-missing indicator $R_{i}^{C}\sim\mathcal{B}(\textup{expit}\{\textup{logit}(1-p_{m})+(C_{i}-N_{i}/10)/2\})$, and treatment $A_{i}\sim\mathcal{B}(0.5)$, where $\mathcal{B}(p)$ represents a Bernoulli distribution with success probability $p$. We then independently generate two individual covariates $X_{i1j}\sim\mathcal{B}(N_{i}/100)$ and $X_{i2j}=b_{i2j}+I\{C_{i}>0\}c_{i}$, where $b_{i2j}\sim N(C_{i}\overline{X}_{i1},1)$ and $c_{i}\sim N(0,1)$ with $\overline{X}_{i1}=\sum_{j=1}^{N_{i}}X_{i1j}/N_{i}$. Their non-missing indicators are generated by $R_{i1j}^{X}\sim\mathcal{B}(1-p_{m})$ and $R_{i2j}^{X}\sim\mathcal{B}(\textup{expit}\{\textup{logit}(1-p_{m})+(C_{i}-N_{i}/10)/2\})$, respectively. Next, we generate the outcome and its missingness indicator from
$Y_{ij}$ $=0.1(R_{i}^{C}C_{i}-1)e^{R_{i1j}^{X}X_{i1j}}|R_{i2j}^{X}(X_{i2j}+1)|+10R_{i1j}^{X}X_{i1j}A_{i}+\gamma_{i}+\varepsilon_{ij},$
$R_{ij}^{Y}$ $\sim\mathcal{B}(\textup{expit}\{\textup{logit}(0.99-0.2p_{m})-(1.5+5p_{m})R_{i1j}^{X}X_{i1j}\}),$
where $\gamma_{i}\sim N(0,1)$ is a cluster-level random intercept and $\varepsilon_{ij}\sim N(0,1)$ is the random error. When we further consider uniform sampling with missing $N_{i}$, we randomly sample $M_{i}$ from integers between $N_{i}/2-3$ and $N_{i}/2+2$ and then uniformly sample $M_{i}$ individuals for cluster $i$, which label as $E_{ij}=1$. Then, the observed data for each cluster $i$ become $\mathbf{O}_{i}=\{(R_{ij}^{Y},R_{ij}^{Y}Y_{ij},A_{i},R_{i1j},R_{i1j}X_{i1j},R_{i2j},R_{i2j}X_{i2j},R_{i}^{C},R_{i}^{C}C_{i},R_{i}^{N}N_{i},N_{i},M_{i}):j=1,\dots,N_{i};E_{ij}=1\}$. Across all eight scenarios, the covariate missingness mechanism includes both missing completely at random and missing not at random, while outcomes follow the rMAR assumption. We generate 1000 data replicates under each scenario to evaluate the performance of the estimators.
We compute five estimators for each data replicate. The unadjusted estimator simply takes the mean difference of cluster-average observed outcomes between treated and control clusters, and its variance is computed by the sandwich variance estimator. The IPW estimator and DR-Aug-GEE estimator are computed using the “CRTgeeDR” package (Prague et al.,, 2017): both estimators use a generalized linear model for outcome missingness, while the latter further involves an outcome linear regression model on main terms of covariates; an independence correlation structure is used for both estimators. For our proposed method, we adopt constant weights and implement both parametric models and machine learning methods for nuisance function estimation, and the resulting estimators are referred to as DR-PM and DR-ML, respectively. DR-PM uses generalized linear models to model treatment assignment and outcome missingness, and adopts generalized estimating equations with an exchangeable working correlation structure for outcome modeling. DR-ML adopts SuperLearner (van der Laan et al.,, 2007) to fit an ensemble model of generalized linear models, random forests, and neural networks for outcome missingness and observed outcomes, but does not model treatment assignment. Across all estimators, the outcome missing model is correctly specified, while the outcome distribution is complex enough such that linear models on main terms of covariates are misspecified. For sandwich variance estimation, we additionally multiply the variance estimator by a factor $m/(m-7)$, where $7$ is the number of covariates, motivated by a straightforward approach in the regression context to correct for small-sample bias (MacKinnon and White,, 1985). This choice improves the coverage probability by 2-5% when $m=30$ and less than 1% when $m=100$. For each estimator, we report their bias for the average treatment effect ($\Delta=5(1-p_{m})$), empirical standard error, average of standard error estimators, and coverage probability of 95% confidence interval based on $t$-distribution with degree of freedom $m-7$.
In the second simulation study, we generate the observed data and compute the DR-PM and DR-ML estimators with variance estimators following the same way as in the first simulation study. Then, we use Equation (5) to search the tipping points such that the proposed estimator is no longer statistically significant at the 0.05 level. For demonstration purposes, we only search for constant tipping points, which assumes $\delta_{a}(M_{i},N_{i})\equiv\delta_{a}$ and $\gamma_{a}(\boldsymbol{B}_{ij})\equiv\gamma_{a}$ are not random variables. Since the data-generating distribution of $R_{ij}^{Y}$ does not depend on $A_{i}$, we can write out the bias of $\widehat{\Delta}^{\dagger}$ for estimating $\Delta$ as
$S(\delta_{1}-\delta_{0},\gamma_{1}-\gamma_{0})=E[(N_{i}-M_{i})/N_{i}](\delta_{1}-\delta_{0})+E[E[1-R_{ij}^{Y}|E_{ij}=1,M_{i},N_{i}]](\gamma_{1}-\gamma_{0}),$
where $E[(N_{i}-M_{i})/N_{i}]$ and $E[E[1-R_{ij}^{Y}|E_{ij}=1,M_{i},N_{i}]]$ can be estimated from data. Finally, we consider $\delta_{1}-\delta_{0}=0,1,2,3,4$, and, for each $\delta_{1}-\delta_{0}$, we search the minimum value of $\gamma_{1}-\gamma_{0}$ such that the bias-corrected 95% confidence interval covers zero, yielding the combinations of $(\delta_{1}-\delta_{0},\gamma_{1}-\gamma_{0})$ as the tipping point. For the 1000 data replicates, we report the mean and standard error of estimated tipping points to reflect the sensitivity of the proposed estimators.
6.2 Simulation results
Table 1 summarizes the results of the first simulation study. Across all eight scenarios, the proposed DR-PM and DR-ML estimators have negligible bias and nominal coverage probabilities, which confirms our theoretical results. Comparing DR-PM and DR-ML, we observe that using machine learning can further reduce the variance by $17\%-36\%$, demonstrating the precision gain from machine learning tools in analyzing CRTs. However, we also notice that this benefit is not for free, that is, machine learning methods lead to $10\%-20\%$ underestimation of variance under our data-generating distribution, causing $1\%-3\%$ under-coverage of confidence intervals.
Unlike the proposed estimators, the three comparison methods all have substantial bias and under-coverage. For the unadjusted estimator, since the missing data are completely ignored, bias presents and becomes larger given more missing data, as expected. While the IPW and DR-Aug-GEE estimators model the missing outcome mechanism, they target a different estimand, $E[N_{i}]^{-1}\{E[N_{i}Y_{ij}|A_{i}=1]-E[N_{i}Y_{ij}|A_{i}=0]\}$ instead of $\Delta$, because they assumed non-informative cluster sizes. This misalignment of estimand leads to bias and inflated type I error rates. Even if the IPW and DR-Aug-GEE estimators target the true estimand, the IPW estimator still has obvious disadvantages in precision ($\sim 10$ fold larger variance than DR-PM) since covariate information is omitted from the outcome model; DR-Aug-GEE performs much better than IPW, but is still 7%-30% less precise than DR-PM because DR-PM further utilizes treatment assignment modeling to improve precision.
Table 1: Results for the first simulation study. ESE: empirical standard error. ASE: average of standard error estimators. CP: coverage probability of 95% confidence intervals based on $t$-distributions.
Sampling $m$ Estimator 10% missing data 30% missing data
Bias ESE ASE CP Bias ESE ASE CP
No 30 Unadjusted -0.27 2.16 2.17 0.95 -1.53 1.32 1.29 0.75
IPW 1.07 2.75 2.58 0.91 0.69 2.11 1.94 0.91
DR-Aug-GEE 0.98 1.12 0.85 0.75 0.71 1.01 0.79 0.83
DR-PM 0.04 1.03 1.00 0.94 -0.08 0.88 0.90 0.95
DR-ML -0.01 0.90 0.84 0.95 -0.02 0.80 0.72 0.92
100 Unadjusted -0.37 1.17 1.19 0.93 -1.50 0.72 0.70 0.44
IPW 1.00 1.50 1.49 0.89 0.76 1.20 1.13 0.88
DR-Aug-GEE 0.99 0.54 0.49 0.47 0.75 0.50 0.46 0.65
DR-PM 0.02 0.50 0.51 0.95 -0.01 0.41 0.44 0.96
DR-ML 0.01 0.40 0.38 0.94 0.00 0.35 0.32 0.94
Yes 30 Unadjusted -0.43 2.27 2.21 0.92 -1.42 1.33 1.31 0.78
IPW 0.90 2.90 2.62 0.88 0.87 2.17 1.99 0.91
DR-Aug-GEE 0.94 1.13 0.88 0.77 0.79 1.10 0.82 0.78
DR-PM 0.05 1.06 1.06 0.94 0.00 0.97 0.98 0.94
DR-ML -0.03 1.00 0.85 0.93 0.00 0.84 0.78 0.93
100 Unadjusted -0.39 1.23 1.19 0.93 -1.50 0.75 0.72 0.45
IPW 0.97 1.58 1.51 0.89 0.76 1.22 1.17 0.89
DR-Aug-GEE 1.00 0.56 0.51 0.52 0.76 0.55 0.49 0.66
DR-PM 0.01 0.54 0.53 0.95 0.00 0.45 0.47 0.96
DR-ML 0.00 0.46 0.42 0.92 0.00 0.39 0.35 0.94
Figure 2 presents the result of the second simulation study. We first consider the special case where $\delta_{1}-\delta_{0}=0$, i.e., no selection bias. From Figure 2, we observe that the proposed methods are less sensitive to missing data assumptions when we have fewer missing data or more clusters. Intuitively, fewer missing data make the unobserved missing data distribution less influential, and more clusters will yield smaller variance, making the boundaries of confidence intervals further away from zero. For the same reason of precision, the DR-ML estimator is also less sensitive than the DR-PM estimator. As $\delta_{1}-\delta_{0}$ becomes larger, selection bias increasingly contributes to the estimation bias, which leads to stronger sensitivity on the missing data assumptions.
Figure 2: Results for the second simulation study. The x-axis $\delta_{1}-\delta_{0}$ measures the selection bias on the average treatment effect from non-uniform sampling, and the y-axis $\gamma_{1}-\gamma_{0}$ measures the difference in the average treatment effect between observed and missing outcomes. In each plot, the dots with error bias represent the mean value plus/minus the standard error obtained from 1000 simulations.
7 Data application
The Work, Family, and Health Study (Work, Family, and Health Study , 2018, WFHS) is a cluster-randomized trial for improving the well-being of employees and reducing their work-family conflict. In this trial, 56 working groups (clusters) in an anonymized company were randomized in a 1:1 ratio to receive a workplace intervention (treatment) or not (control). Each cluster has 7-60 employees (mean 24, standard error 12), while participants of this CRT range 3-42 per cluster (mean 15, standard error 8). The outcome we focus on is control over work hours defined at the 6-month follow-up, a continuous measure ranging from 1 to 5. We include two covariates: the baseline value of the outcome measure as the individual level covariate, and group job functions (core or supporting) at the cluster level. In the data set, 20% of the outcomes are missing, and 13% of the individual-level covariates are missing, while the cluster-level covariate has no missingness. This trial adopted a biased-coin adaptive design for treatment allocation, but for demonstration purposes, we assume simple randomization; this simplification will lead to slight overestimation of the variance but has no impact on the estimates, and we will return to a discussion of this topic in Section 8. We consider estimating the average treatment effect defined in (1) using the five estimators defined in Section 6.
Table 2 summarizes the results of our data application. Across all five estimators, their point estimates are relatively close, and their confidence intervals all exclude zero. Since we are unaware of the truth and have no data replications, the similarities among estimators can result from a missing-completely-at-random data-generating distribution or simply randomness. In terms of precision, covariate adjustment provides substantial precision gain, and machine learning remains a valid tool for further reducing variance (36%) compared to parametric models. Of note, while we observe that DR-Aug-GEE leads to a similar variance to DR-ML, this is because DR-Aug-GEE targets a different estimand (as discussed in Section 3.1), for which the associated asymptotic variance can be systematically smaller even given the same nuisance function estimators (e.g., comparing DR-Aug-GEE and DR-PM), resulting in incomparable variance estimators.
Table 2: Results of re-analysis of the WFHS data.
Estimate Standard error 95% confidence interval
Unadj 0.17 0.09 (0.00, 0.34)
IPW 0.18 0.07 (0.03, 0.32)
DR-Aug-GEE 0.22 0.05 (0.12, 0.32)
DR-PM 0.20 0.06 (0.07, 0.32)
DR-ML 0.20 0.05 (0.10, 0.30)
In addition, we performed a sensitivity analysis for our proposed estimators to violations of missing data assumptions. Like the simulation studies, we adopt constant $\delta_{1}-\delta_{0}$ and $\gamma_{1}-\gamma_{0}$. If there is no selection bias, i.e., $\delta_{1}-\delta_{0}=0$, the tipping point for DR-ML is 0.74. It means that our conclusion that the intervention is effective remains unchanged as long as the treatment effect for individuals with missing outcomes is no smaller than $-0.54$. Since the outcome measures range 1-5, a treatment effect of $-0.54$ can be regarded as a substantial negative effect, which in turn implies that our conclusion based on DR-ML is relatively insensitive to missing not at random. If the outcome is missing at random, i.e., $\gamma_{1}-\gamma_{0}=0$, the tipping point is 0.28 for DR-ML, meaning that we can tolerate a $-0.08$ treatment effect for the non-participants. Compared to DR-ML, the tipping points for DR-PM are 0.35 and 0.21 for the above two scenarios, representing a weakened insensitivity due to the increased variance.
8 Discussion
As CRTs become increasingly popular for pragmatic clinical trials evaluating interventions in routine practice conditions, our contribution to improving the analysis of CRTs is three-fold. First, under clearly listed assumptions, we propose a doubly-robust estimator that can simultaneously handle many commonly seen missing data types. Second, we propose three techniques to further improve the precision, aiming to improve the study power and accelerate CRTs to generate real-world evidence. Third, we propose a practical sensitivity analysis to evaluate the influence of violations on missing data assumptions. Taken together, our results not only broaden the development of statistical methodology for CRTs, but are also informative to the design and analysis of future CRTs.
While we focus on simple randomization (i.e., assigning treatment via independent coin flips) in this article, covariate-adaptive randomization is also frequently used in practice, aiming to achieve treatment balance within each covariate stratum of interest (e.g., hospital site, age group). Examples of covariate-adaptive randomization for CRTs include stratified block randomization, biased-coin randomization, and rerandomization. For these restricted designs, Lu et al., (2023); Wang et al., (2023); Rafi, (2023) have developed the asymptotic theories, which all showed that covariate-adaptive randomization retains the validity and can lead to a smaller variance for a wide range of estimators including our proposed ones. Given these results, deriving the analytical formula to account for this variance reduction can be a useful future extension.
Given within-cluster sampling, we characterize a basic uniform-sampling scenario where enrollment causes no selection bias. If selection bias may occur, principal stratification (Li et al.,, 2022; Papadogeorgou et al.,, 2023) with structural assumptions can be used to identify treatment effects for certain subpopulations, e.g., individuals that are always enrolled regardless of treatment assignment. However, how the principal stratification method can accommodate missing outcomes and covariates remains an open question.